# Symmetry-enriched topological order and quasi-fractonic behavior in $\mathbb{Z}_{N}$ stabilizer codes

### v1

@@ V.1 Quasi-fractonic behavior of BB codes
Figure 3: The quasi-fractonic behavior of the DCY model. Here $N=4$ and $a=1$. This figure depicts the process of moving excitations in the DCY model. The gray dots represent $Z$-type stabilizers at each position, where we only draw the stabilizers along the $x$-directions in each step. The yellow dots represent the excitations, denoted by the monomials next to the dots. In Step 1, the anyonic excitations are split by moving. After Step 3, the original excitation is moved from $(0,0)$ to $(4,0)$.
@@
Like the binary BB code [50], the anyons of $\mathbb{Z}_{N}$ qudit BB codes cannot be moved smoothly via local operators in general, which we refer to as the quasi-fractonic behavior of BB codes.
@@
To illustrate the quasi-fractonic behavior of $\mathbb{Z}_{N}$ BB codes, we consider a concrete example: the DCY model, assuming for simplicity that $a$ and $N$ are coprime. We consider the $Z$-type excitations, which can be labeled as an element in $R$. Initially, assume that the lowest energy excitation is located at the origin, denoted as $1$. Annihilating this excitation by performing a qudit operation $X_{(2,{1}/{2})}^{-1}$ will create two new excitations: one located at $(1,0)$ and the other located at $(2,0)$, and these excitations are collectively denoted by $(a+1)x-x^{2}$. This indicates that moving anyons by performing local operations may cause them to split, in contrast to the toric code, where they can be moved freely without creating additional excitations. This process is depicted in Step 1 in Fig. 3.
@@
These behaviors are quite similar to the fractons, whose excitations are immobile. However, unlike the fractons, anyon excitations remain mobile for particular distances. Hence, we refer to these exotic behaviors as quasi-fractonic behaviors. For example, we consider the DCY model with $N=4$ and $a=1$. In this case, the above process splits the original excitation into two anyons denoted as $3x^{2}+2x$. However, as depicted in Fig. 3(a), performing operations $X_{(3,{1}/{2})}^{2}$ and $X_{(4,{1}/{2})}$ successively annihilates all other excitations and creates a new excitation $x^{4}$, indicating that the original anyon is moved from $(0,0)$ to $(4,0)$. For a comprehensive approach to the mobility of anyons in the DCY model, see Sec. VI.2.
@@ V.2 SET orders and mobility
The quasi-fractonic behavior described above is rooted in the symmetry-enriched topological (SET) order feature, where the translation symmetry acts nontrivially on anyons and permutes anyon types [45, 50]. To see this, we return to the DCY model where $a$ and $N$ are coprime. The anyon types of its $Z$-type excitations are labeled by the basis $\{1,x,y,xy\}$ by long divisions of polynomials. A $Z$-type excitation represented by $1$, and its translation represented by $x$, belong to different anyon types. Consequently, moving an anyon between these two positions via local operations is forbidden.
However, according to Corollary 1, the total number of anyon types of topological BB codes is finite. This implies that the anyon type of the initial excitation resurrects after finite steps of translation actions. Thus, the anyons are still mobile with particular distances for qudit BB codes, distinguishing them from fractons.
For a general $\mathbf{BB}_{N}(f,g)$ code, its SET feature and mobility are characterized by its mobility group. The mobility group is a subgroup of the translation group $\Lambda$, and is explicitly defined as:
$\Lambda_{\mathcal{C}}$ $=\{\lambda\in\Lambda|\lambda\theta=\theta,\theta\in\mathcal{C}\}$
$=\{\lambda\in\Lambda|\lambda-1\in(f,g)_{R}\}.$ (27)
The mobility group describes all translation actions preserving anyon types, and thus identifies all possible displacements of anyons by performing local operations on the lattice.
In particular, we define the anyon periodicities along the $x$- or $y$-directions as the minimum distance after which anyon types reappear under translation. These periodicities, denoted as $l_{x}$ and $l_{y}$, are given by
$l_{x}$ $=\min\{\alpha|x^{\alpha}-1\in(f,g)_{R}\},$ (28)
$l_{y}$ $=\min\{\beta|y^{\beta}-1\in(f,g)_{R}\},$
For product codes, the polynomials specifying $f$ and $g$ become univariate. In this case, the mobility group $\Lambda_{C}$ is characterized by its periodicities:
@@
where $\langle x^{l_{x}},y^{l_{y}}\rangle$ denote the subgroup of the translation group generated by $x^{l_{x}}$ and $y^{l_{y}}\in\Lambda$.
@@ V.3 Ground state degeneracy
Another feature of translation symmetry enriched topological orders, absent in conventional topological orders, is the possible dependence of ground-state degeneracy (GSD) on the system size [53, 50]. In this section, we describe how to evaluate the GSD of the BB code on a torus of size $L_{x}\times L_{y}$ and discuss its relation to the underlying SET order in detail.
@@
$R_{L}\xrightarrow{\partial_{L}=\left(\begin{array}[]{cc}\partial_{AL}&0\\ 0&\partial_{BL}\end{array}\right)}R_{L}^{2}\xrightarrow{\varepsilon=\left(\begin{array}[]{cc}0&\varepsilon_{AL}\\ \varepsilon_{BL}&0\end{array}\right)}R_{L},$ (31)
@@
Here, $S=\langle S_{A},S_{B}\rangle$ is the stabilizer group, where $S_{A}$ and $S_{B}$ are the groups generated by the operators $\{A_{(i,j)}\}$ and $\{B_{(i,j)}\}$ (defined in Eq. (2)) over all sites. The evaluation of this trace yields:
@@
Clearly, the GSD depends on the system’s size. According to Eq. (23) and Eq. (33), it cannot exceed its total number of anyon types, and reaches this maximum value if and only if the system size is an integer multiple of the periodicity given in Eq. (28).
This reflects the SET nature of BB codes: The translation symmetry acts non-trivially on the anyons, so unless the system’s size matches anyon periodicities, not all anyon types are preserved when they traverse non-contractible loops on a torus. Consequently, non-contractible loop operators do not exist for some anyon types, which in turn reduces the GSD.
@@ VI Analytical results for explicit models
This section analytically characterizes the essential topological properties of the models introduced in Sec. II. First, in Sec. VI.1, we simply and concisely reproduce the known topological properties of the WCF model [53]. Next, in Sec. VI.2, our analysis of the DCY model resolves its mobility puzzle and provides what we believe to be the first accurate determination of its ground-state degeneracy (GSD). Finally, in Sec. VI.3 and Sec. VI.4, we apply Theorem 2 to determine the fusion rules for the product and BB codes with the toric layout. These results demonstrate the broad applicability of our framework for characterizing topological order.
@@ VI.1 WCF model
Anyon types and fusion rules—We specify the anyon type of the WCF model by invoking Theorem 2. For each prime factor $p_{i}$ of the qudit dimension $N=\prod_{i=1}^{n}p_{i}^{k_{i}}$, the corresponding $\mathbb{Z}_{p_{i}}$ WCF model is in the trivial phase if $p_{i}$ divides $a$, but has fusion rules $\mathbb{Z}_{p_{i}}^{2}$ if $a$ and $p_{i}$ are coprime. Thus, the fusion rules of the WCF model are specified by
@@
Mobility—According to Eq. (28), the anyon types remain unchanged under translations by $\alpha$ and $\beta$ along $x$- and $y$-directions exactly when
@@
where $\alpha,\beta\in\mathbb{Z}$. This gives the mobility conditions of anyons. The anyon periodicities $l_{x}$ and $l_{y}$ are given by the smallest positive $\alpha$ and $\beta$ satisfying Eq. (36a) and Eq. (36b), respectively.
We focus first on the $x$-direction. We decompose the base ring $\mathbb{Z}_{N}=\mathbb{Z}_{\frac{N}{N_{a}}}\oplus\mathbb{Z}_{N_{a}}$ using the Chinese remainder theorem. It is clear that the following equalities hold:
@@
Thus, the mobility condition in Eq. (36a) becomes $a^{\alpha}=1\bmod N_{a}$ since $x^{\alpha}-1$ needs to vanish on both direct summands. Euler’s theorem guarantees the existence of such an $\alpha$. We denote the smallest positive integer $\alpha$ satisfying $a^{\alpha}=1\bmod N_{a}$ by $\operatorname{ord}_{N_{a}}(a)$. It is also the order of $a$ in the multiplicative group of units of $\mathbb{Z}_{N_{a}}$. Thus, the periodicity along the $x$-direction is $\operatorname{ord}_{N_{a}}\left(a\right)$.
The periodicity along the $y$-direction can be obtained analogously. We conclude that the periodicities of the WCF model along the $x$- and $y$-directions are
@@
Ground state degeneracy—The GSD of the WCF model equals $|R_{L}/(f,g)|^{2}$ according to Eq. (33). By decomposing $R_{L}$ into a direct sum, we get
@@ VI.2 DCY model
Anyon types and fusion rules—In analogy with the WCF model, applying Theorem 2, the fusion rules of the DCY model are given by
@@
Mobility—Similar to the WCF model, the periodicities $l_{x}$ and $l_{y}$ are the smallest positive $\alpha$ and $\beta$ satisfying the mobility conditions in Eq. (36a) and Eq. (36b). We begin by concentrating on the $x$-direction. It is evident that
@@
denotes the finite geometric series in the variable $a$, up to degree $\alpha-1$. We decompose the base ring into $\mathbb{Z}_{N}=\mathbb{Z}_{\frac{N}{N_{a}}}\oplus\mathbb{Z}_{N_{a}}$. According to Eq. (44a) and Eq. (44b), the mobility condition in Eq. (36a) is trivial on the first direct summand, and reduces to $G_{\alpha}=0\bmod N_{a}$ on the second. This condition implies $a^{\alpha}=1\bmod N_{a}$. Therefore, the periodicity $l_{x}$ must be an integer multiple of $\operatorname{ord}_{N_{a}}(a)$. Additionally, using the property $a^{m+\operatorname{ord}_{N_{a}}(a)}=a^{m}\bmod N_{a}$, it can be shown that $nG_{\operatorname{ord}_{N_{a}}(a)}=G_{n\operatorname{ord}_{N_{a}}(a)}\bmod N_{a}$ for any positive integer $n$. Thus, basic algebraic manipulation gives
@@
Therefore, the periodicities of the DCY model along $x$- and $y$-directions are
@@
All anyons in the DCY model are able to hop between lattice sites separated by displacements belonging to $\Lambda_{{\operatorname{DCY}}}$. This resolves the anyon mobility puzzle of the DCY model.
Ground state degeneracy—The GSD of the DCY model is given by $|R_{L}/(f,g)|^{2}$ with $f,g$ satisfying Eq. (20). In this case, the quotient module $R_{L}/(f,g)$ still decomposes according to the direct sum in Eq. (40), so we analyze the two summands separately.

### v2

@@ V.1 Quasifractonic mobility of anyons
Figure 3: The quasifractonic behavior of the DCY model. Here $N=4$ and $a=1$. This figure depicts the process of moving excitations in the DCY model. The gray dots represent $Z$-type stabilizers at each position, where we only draw the stabilizers along the $\hat{x}$ directions in each step. The yellow dots represent the excitations, denoted by the monomials next to the dots. In step 1, the anyonic excitations are split by moving. After step 3, the original excitation is moved from $(0,0)$ to $(4,0)$.
@@
Like the qubit BB codes [50], anyons in $\mathbb{Z}_{N}$ BB codes typically cannot be moved smoothly by local operators.
To illustrate this, we consider a concrete example: the DCY model, and for simplicity assume that $a$ and $N$ are coprime [see Eq. (20)]. A complete treatment of anyon mobility in this model is presented in Sec. VI.2.
@@
We focus on $Z$-type excitations, which can be labeled as an element in $R$. Initially, assume that the lowest energy excitation is located at the origin, denoted as $1$. Annihilating this excitation by performing a qudit operation $X_{(2,{1}/{2})}^{-1}$ will create two new excitations: one located at $(1,0)$ and the other located at $(2,0)$, and these excitations are collectively denoted by $(a+1)x-ax^{2}$. This indicates that moving anyons by performing local operations may cause them to split, in contrast to the toric code, where they can be moved freely without creating additional excitations. This process is depicted in step 1 in Fig. 3.
@@
This behavior resembles that of fractons, which split into multiple pieces when one attempts to move them. However, unlike the fractons, anyon excitations still permit one-to-one hopping over certain finite distances. We refer to this behavior as quasifractonic mobility. As an explicit example, consider the DCY model with $N=4$ and $a=1$. In this case, the above process splits the original excitation into two anyons denoted as $3x^{2}+2x$. However, as depicted in Fig. 3, performing operations $X_{(3,{1}/{2})}^{2}$ and $X_{(4,{1}/{2})}$ successively annihilates all other excitations and creates a new excitation $x^{4}$, indicating that the original anyon is moved from $(0,0)$ to $(4,0)$.
@@ V.2 SET orders and mobility
A central question arises: Is it universally true for topological qudit BB codes that their topological excitations—though fractonlike under local moves—always permit one-to-one hopping over sufficiently large but finite distances? We answer this question in the affirmative below. Apparent discrepancies in the literature (e.g., the DCY model with non-coprime parameters $N=12$ and $a=2$ [54]) often stem from computational oversights in specific cases.
A general resolution of this mobility puzzle is naturally formulated in terms of SET order, which captures the interplay between symmetry and topological order. Here, the relevant symmetry is lattice translation, which is often essential for understanding fracton and fracton-like phenomena [102, 103, 104]. The obstruction to elementary hopping [e.g., $(0,0)\to(1,0)$] arises because translations act nontrivially on anyons by permuting anyon types [99, 50]. Generically, moving an anyon $\eta\in\mathcal{C}$ by a lattice vector $(i,j)$ is forbidden if and only if $\eta$ and its translate $x^{i}y^{j}\eta$ represent distinct anyon types.
Thus, to identify all lattice displacements that admit one-to-one anyon transport, we introduce the mobility group $\Lambda_{\mathcal{C}}$. It is a subgroup of the translation group $\Lambda$, defined by
$\Lambda_{\mathcal{C}}$ $=\{\lambda\in\Lambda|\lambda\eta=\eta,\forall\eta\in\mathcal{C}\}$
$=\{\lambda\in\Lambda|\lambda-1\in(f,g)_{R}\},$ (27)
where $(f,g)_{R}$ denotes the ideal of $R$ generated by polynomials $f$ and $g$. Physically, $\Lambda_{\mathcal{C}}$ consists of precisely those translations that leave all anyon types invariant; consequently, it specifies exactly which lattice displacements admit one-to-one anyon transport implemented by string operators.
As a key aspect of $\Lambda_{\mathcal{C}}$, we define the anyon periodicities along the $\hat{x}$ and $\hat{y}$ directions as the minimal translation distances under which all anyon types remain invariant. These periodicities, denoted by $l_{x}$ and $l_{y}$, are given by
$l_{x}$ $=\min\{\alpha\in\mathbb{Z}_{>0}\mid x^{\alpha}-1\in(f,g)_{R}\},$ (28)
$l_{y}$ $=\min\{\beta\in\mathbb{Z}_{>0}\mid y^{\beta}-1\in(f,g)_{R}\}.$
Crucially, $l_{x}$ and $l_{y}$ always exist and are finite for any topological BB code, regardless of whether it is a qubit or qudit system. This follows from Corollary 1, which establishes that the set of anyon types $\mathcal{C}$ is finite. Consequently, $\mathcal{C}$ forms a finite $\Lambda$-set under lattice translations, ensuring that every translation orbit must close. This guarantees the existence of minimal positive integers $l_{x}$ and $l_{y}$ satisfying Eq. (28). This finiteness ensures that all excitations in topological BB codes—while appearing fracton-like at length scales below $l_{x}$ and $l_{y}$—possess long-range one-to-one mobility, fundamentally distinguishing them from true fractons.
For HGP codes, $l_{x}$ and $l_{y}$ completely characterize $\Lambda_{\mathcal{C}}$. In this case, the polynomials specifying $f$ and $g$ become univariate. As a result, the mobility group $\Lambda_{C}$ takes the form
@@
where $\langle x^{l_{x}},y^{l_{y}}\rangle$ denotes the subgroup of the translation group generated by $x^{l_{x}}$ and $y^{l_{y}}\in\Lambda$. For general BB codes with multivariate defining polynomials, $\langle x^{l_{x}},y^{l_{y}}\rangle$ may only be a subgroup of $\Lambda_{\mathcal{C}}$.
We conclude this subsection by examining the DCY model with parameters $N=12$ and $a=2$ as an explicit illustration of the general framework. This specific case is of particular interest because the existing literature [54] has questioned the finiteness of its periodicities, $l_{x}$ and $l_{y}$. For this model, the defining polynomials are f=(x1)(x2)f=(x-1)(x-2) and g=(y1)(y2)g=(y-1)(y-2), as given in Eq. (20). To analyze the mobility, it suffices to consider the charge sector $\mathcal{Q}$ of the anyon module $\mathcal{C}=\mathcal{Q}\oplus\overline{\mathcal{Q}}$:
$\mathcal{Q}=\frac{\mathbb{Z}_{12}[x^{\pm},y^{\pm}]}{(f,g)}\cong\frac{\mathbb{Z}_{3}[x^{\pm},y^{\pm}]}{(f,g)}\oplus\frac{\mathbb{Z}_{4}[x^{\pm},y^{\pm}]}{(f,g)},$ (30)
where we utilize the decomposition $\mathbb{Z}_{12}\cong\mathbb{Z}_{3}\oplus\mathbb{Z}_{4}$. We note
  1. 1.
    In the $\mathbb{Z}_{3}$ sector, $2\equiv-1\pmod{3}$, so $f=(x-1)(x+1)=x^{2}-1$ and $g=y^{2}-1$.
  2. 2.
    In the $\mathbb{Z}_{4}$ sector, (x2)(x-2) is invertible since $(x-2)(x+2)x^{-2}\equiv 1\pmod{4}$. Similarly, (y2)(y-2) is also invertible.
Consequently, the charge module simplifies to:
$\mathcal{Q}\cong\frac{\mathbb{Z}_{3}[x^{\pm},y^{\pm}]}{(x^{2}-1,y^{2}-1)}\oplus\frac{\mathbb{Z}_{4}[x^{\pm},y^{\pm}]}{(x-1,y-1)}.$ (31)
Under this direct sum decomposition of $\mathcal{Q}$, a unit $X$-type excitation at the origin is represented by $1\choose 1$. Translation along $\hat{x}$ acts via multiplication by $x$, yielding the orbit (in the quotient)
${1\choose 1}\xrightarrow{\cdot x}{x\choose x}\equiv{x\choose 1}\xrightarrow{\cdot x}{x^{2}\choose x^{2}}\equiv{1\choose 1}$. (32)
This confirms $l_{x}=2$, contradicting the claim of non-existence of finite periods in Ref. [54]. A quick check confirms this: The $Z$-string operator $Z_{(\frac{1}{2},0)}Z_{(-\frac{1}{2},0)}^{3}Z_{(-\frac{3}{2},0)}^{6}$ successfully moves a unit $X$-type excitation from $(0,0)$ to $(2,0)$. Section VI.2 extends this analysis to arbitrary DCY models, further confirming the general result on the existence of one-to-one anyon hopping processes.
@@ V.3 Ground state degeneracy
Another feature of translation symmetry enriched topological orders, absent in conventional topological orders, is the possible dependence of GSD on the system size [53, 50]. In this section, we describe how to evaluate the GSD of the BB code on a torus of size $L_{x}\times L_{y}$ and discuss its relation to the underlying SET order in detail.
@@
$R_{L}^{2}\xrightarrow{\partial_{L}=\left(\begin{array}[]{cc}\partial_{AL}&0\\ 0&\partial_{BL}\end{array}\right)}R_{L}^{4}\xrightarrow{\varepsilon_{L}=\left(\begin{array}[]{cc}0&\varepsilon_{AL}\\ \varepsilon_{BL}&0\end{array}\right)}R_{L}^{2},$ (34)
@@
Here, $S=\langle S_{A},S_{B}\rangle$ is the stabilizer group, where $S_{A}$ and $S_{B}$ are the groups generated by the operators $\{A_{(i,j)}\}$ and $\{B_{(i,j)}\}$ [defined in Eq. (2)] over all sites. The evaluation of this trace yields:
@@
Clearly, the GSD depends on the system’s size. According to Eq. (23) and (36), it cannot exceed its total number of anyon types, and reaches this maximum value if and only if the system size is an integer multiple of the periodicity given in Eq. (28).
This reflects the SET nature of BB codes: The translation symmetry acts nontrivially on the anyons, so unless the system’s size matches anyon periodicities, not all anyon types are preserved when they traverse non-contractible loops on a torus. Consequently, non-contractible loop operators do not exist for some anyon types, which in turn reduces the GSD.
@@ VI Analytical results for explicit models
This section analytically characterizes the essential topological properties of the models introduced in Sec. II. First, in Sec. VI.1, we simply and concisely reproduce the known topological properties of the WCF model [53]. Next, in Sec. VI.2, our analysis of the DCY model resolves its mobility puzzle and provides what we believe to be an accurate determination of its GSD. Finally, in Secs. VI.3 and Sec. VI.4, we apply Theorem 2 to determine the fusion rules for the product and BB codes with the toric layout. These results demonstrate the broad applicability of our framework for characterizing topological order.
@@ VI.1 WCF model
VI.1.1 Anyon types and fusion rules
We specify the anyon type of the WCF model by invoking Theorem 2. For each prime factor $p_{i}$ of the qudit dimension $N=\prod_{i=1}^{n}p_{i}^{k_{i}}$, the corresponding $\mathbb{Z}_{p_{i}}$ WCF model is in the trivial phase if $p_{i}$ divides $a$, but has fusion rules $\mathbb{Z}_{p_{i}}^{2}$ if $a$ and $p_{i}$ are coprime. Thus, the fusion rules of the WCF model are specified by
@@ VI.1.1 Anyon types and fusion rules
VI.1.2 Mobility
According to Eq. (28), the anyon types remain unchanged under translations by $\alpha$ and $\beta$ along $\hat{x}$ and $\hat{y}$ directions exactly when
@@ VI.1.2 Mobility
where $\alpha,\beta\in\mathbb{Z}$. This gives the mobility conditions of anyons. The anyon periodicities $l_{x}$ and $l_{y}$ are given by the smallest positive $\alpha$ and $\beta$ satisfying Eqs. (39a) and (39b), respectively.
We focus first on the $\hat{x}$ direction. We decompose the base ring $\mathbb{Z}_{N}=\mathbb{Z}_{\frac{N}{N_{a}}}\oplus\mathbb{Z}_{N_{a}}$ using the Chinese remainder theorem. It is clear that the following equalities hold:
@@
Thus, the mobility condition in Eq. (39a) becomes $a^{\alpha}=1\bmod N_{a}$ since $x^{\alpha}-1$ needs to vanish on both direct summands. Euler’s theorem guarantees the existence of such an $\alpha$. We denote the smallest positive integer $\alpha$ satisfying $a^{\alpha}=1\bmod N_{a}$ by $\operatorname{ord}_{N_{a}}(a)$. It is also the order of $a$ in the multiplicative group of units of $\mathbb{Z}_{N_{a}}$. Thus, the periodicity along the $\hat{x}$ direction is $\operatorname{ord}_{N_{a}}\left(a\right)$.
The periodicity along the $\hat{y}$ direction can be obtained analogously. We conclude that the periodicities of the WCF model along the $\hat{x}$ and $\hat{y}$ directions are
@@
VI.1.3 Ground state degeneracy
The GSD of the WCF model equals $|R_{L}/(f,g)|^{2}$ according to Eq. (36). By decomposing $R_{L}$ into a direct sum, we get
@@ VI.2 DCY model
VI.2.1 Anyon types and fusion rules
—In analogy with the WCF model, applying Theorem 2, the fusion rules of the DCY model are given by
@@ VI.2.1 Anyon types and fusion rules
VI.2.2 Mobility
Similar to the WCF model, the periodicities $l_{x}$ and $l_{y}$ are the smallest positive $\alpha$ and $\beta$ satisfying the mobility conditions in Eqs. (39a) and (39b). We begin by concentrating on the $\hat{x}$ direction. It is evident that
@@ VI.2.2 Mobility
denotes the finite geometric series in the variable $a$, up to degree $\alpha-1$. We decompose the base ring into $\mathbb{Z}_{N}=\mathbb{Z}_{\frac{N}{N_{a}}}\oplus\mathbb{Z}_{N_{a}}$. According to Eqs. (47a) and (47b), the mobility condition in Eq. (39a) is trivial on the first direct summand, and reduces to $G_{\alpha}=0\bmod N_{a}$ on the second. This condition implies $a^{\alpha}=1\bmod N_{a}$. Therefore, the periodicity $l_{x}$ must be an integer multiple of $\operatorname{ord}_{N_{a}}(a)$. Additionally, using the property $a^{m+\operatorname{ord}_{N_{a}}(a)}=a^{m}\bmod N_{a}$, it can be shown that $nG_{\operatorname{ord}_{N_{a}}(a)}=G_{n\operatorname{ord}_{N_{a}}(a)}\bmod N_{a}$ for any positive integer $n$. Thus, basic algebraic manipulation gives
@@
Therefore, the periodicities of the DCY model along $\hat{x}$ and $\hat{y}$ directions are
@@
The mobility group $\Lambda_{{\operatorname{DCY}}}$ gives a precise characterization of the DCY model’s anyon mobility and resolves the anyon mobility puzzle of the DCY model. It never becomes trivial, so the one-to-one hoppings are always permitted, aligning with the argument in Sec. V.2.
VI.2.3 Ground state degeneracy
The GSD of the DCY model is given by $|R_{L}/(f,g)|^{2}$ with $f,g$ satisfying Eq. (20). In this case, the quotient module $R_{L}/(f,g)$ still decomposes according to the direct sum in Eq. (43), so we analyze the two summands separately.